OpenStudy (anonymous):

Simplify the expression below. –(–7y + 12)

OpenStudy (anonymous):

Do you know how to distribute?

OpenStudy (anonymous):

7y-12

OpenStudy (anonymous):

7y-12

Parth (parthkohli):

-1(-7y + 12) = 7y + ( - 12) = 7y - 12

OpenStudy (anonymous):

The reason why it's 7y-12 is because they distribute the negative sign. Which essentially means multiply everything by the negative and when you do that it just switches the sign.

OpenStudy (anonymous):

first distribute, multiplication of the same signs is always positive (like (-)(-) and (+)(+)) and multiplication of the opposite is (like (-)(+) or vice versa). so : –(–7y + 12) you have this: (–1)(–7y + 12) which is " (–1)(–7y ) + (-1)(+12) (-)(-) = + and (-)(+) = - 7y - 12

OpenStudy (anonymous):

multiplication of the same signs is always positive (like (-)(-) and (+)(+)) and multiplication of the opposite is always negative (like (-)(+) or vice versa). :))